OpenStudy (anonymous):

The formula r=2√5L can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length L, in feet. How far will a car skid at 30mph? The car will skid ? feet.

OpenStudy (anonymous):

30=2sqr(5L) 15=sqr(5L) 15^2=5L L=15^2/5 =45 feet

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

No problems! :)